OpenStudy (anonymous):

Can anyone help me with PreCalc? Write the equation of an exponential function given the points (1,48) and (2,24)?

ganeshie8 (ganeshie8):

hint : 48/2 = 24

OpenStudy (anonymous):

So... in A(t) = A0*e^(-k*t) 1/2 =A? I'm sorry, my teacher has taught me nothing this year, and expects us to know this.

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

forget about e

ganeshie8 (ganeshie8):

:)

ganeshie8 (ganeshie8):

simpler than that

ganeshie8 (ganeshie8):

since going from 1 to 2, is halving the value, the exponential funciton shud be of form :- \(\large y = k \frac{1}{2^x}\)

OpenStudy (anonymous):

Okay... so how would I find the equation given the points?

ganeshie8 (ganeshie8):

since y = 48, when x = 1, k = 96

ganeshie8 (ganeshie8):

so the exponential equation wud be \(\large y = 96 \times 2^{-x}\)

ganeshie8 (ganeshie8):

see if that makes some sense

OpenStudy (anonymous):

Alright, I'll try to figure it out from there(: Thanks!

ganeshie8 (ganeshie8):

good :) u will have to see it for urself i guess